OpenStudy (anonymous):

∭▒〖∇.F ̅.dv〗=∭▒〖F ̅.ds〗 ∇.F ̅ = ∂/∂x (3x+z^77 )+∂/∂y (y^2-sin x^(2 ) z)+∂/∂z (xz+〖ye〗^5x ) ∇.F ̅ = =(3+z^77 )+(2y-sin x^(2 ) z)+(x+〖ye〗^5x ) ∭▒〖∇.F ̅= 〗 ʃ^1 ʃ^3 ʃ^2 (3+z^77 )+(2y-sin x^(2 ) z)+(x+〖ye〗^5x ) .dx dy dz = ʃ^1 ʃ^3 ʃ^2 x+5y . dx dy dz ʃ^1 x.dx = [ x^2/2 ]^1 = [ 1/2 -0 ] = 1/2

OpenStudy (fibonaccichick666):

you my friend need to learn LaTex lol gimme a min to decipher

OpenStudy (anonymous):

I have problem when I typed the question sorry sorry

OpenStudy (fibonaccichick666):

lol it's cool

OpenStudy (fibonaccichick666):

ok so what is the interval or C?

OpenStudy (fibonaccichick666):

\[∭[∇.F̅.dv]=∭[F̅.ds]\] \[∇.F̅=∂/∂x(3x+z77)+∂/∂y(y2−sinx(2)z)+∂/∂z(xz+〖ye〗5x)\] \[∇.F̅==(3+z77)+(2y−sinx(2)z)+(x+〖ye〗5x)\] \[∭〖∇.F̅=〗ʃ1ʃ3ʃ2(3+z77)+(2y−sinx(2)z)+(x+〖ye〗5x).dxdydz\] \[=\int \int \int x+5y.dxdydz\] \[ \int 1x.dx=[x2/2]1=[1/2−0]=1/2\]

OpenStudy (anonymous):

If D = (3x+z^77)ax + (y^2-sin x^2z)ay + (xz +ye^5x)az. Find the electric flux through the surface of the box; 0<x<1 ; 0<y<3 ; 0<z<2

OpenStudy (fibonaccichick666):

so can you tell me what the normal vector is?

OpenStudy (fibonaccichick666):

correction, unit normal vector sorry